Calculate the energy of electron transitions in a one-electron (bohr) system

Question

What is the wavelength of a photon that will excite an electron from n=3 to n=5 in a hydrogen atom?

  • Use R∞=2.179×10−18J for the hydrogen atom Rydberg constant.
  • Use h=6.626×10−34 Js for Planck's constant.
  • Use c=2.998×108ms for the speed of light.
Answer #1

the energy of electron transitions in a one-electron (bohr) system :

wavelength of a photon that will excite an electron from n=3 to n=5 in a hydrogen atom?

We have :

  • R∞=2.179×10−18J ( hydrogen atom Rydberg constant.}
  • h=6.626×10−34 Js ( Planck's constant.)
  • c=2.998×108 ms ( speed of light

from one-electron bohr model dE =  (hc/ ) = R∞ ((1/n12) -(1/n22)

  • we have, n1= 3 (obital from where electron under going excitation by absorption of  light : lower in energy (more negative energy)) ;  

    n2=5 (obital where excited electron will reach : higher in energy (less negative energy).

dE =  (hc/ ) =2.179×10−18J  ((1/32) -(1 / 52))

dE =  (hc/ ) = 1.5495×10−19J

(6.626×10−34 Js* 2.998×108 ms/ ) = 1.5495×10−19J

= 1.282*10−6 m

(1 nm = 10-9 m )

= 1282 nm